how to figure out the missing depreciation expense on an income statement

a. Net sales = $12,150,000.
b. Cost of goods sold = $8,700,000.
c. Other operating expenses = $230,000.
d. Addition to retained earnings = $1,140,000.
e. Dividends paid to preferred and common stockholders = $365,000.
f. Interest expense = $920,000.
The firm’s tax rate is 30 percent.

Calculate the depreciation expense for Corky’s Bedding Corp.
Answer #1

Solution :-

Retained earnings = 1140000

Dividend to preference and equity = 365000

Therefore earning after tax = 1140000 + 365000 = 1505000

Earnings before tax = EAT/(1-Tax) =

= 1505000/(1-0.30) = 2150000

EBIT = EBT + Interest = 2150000 + 920000 = $3070000

Depreciation{Non Cash} = Sales - COGS - Operating expense - EBIT = 12150000 - 8700000 - 230000 - 3070000 = $150000